27. Consider an economy that produces and consumes Bread and Automobile. Data for two different years 2008 and 2011 is given in the following table.
Year 2008 2011
Price of Automobiles $15,000 $20,000
Price of a loaf of bread $5 $8
Number of automobiles produced 200 320
Number of loaves of bread produced 100,000 200,000
Using the year 2008 as a base year,
a. Calculate the nominal and real GDP of 2011.
b. Find the value of GDP Deflator for the year 2011 and interpret.
c. Calculate the inflation rate in 2011.
1
Expert's answer
2020-10-05T14:23:32-0400

"Q_1" Nominal GDP of 2011=price of bread(2011)"\times" quantity of bread (2011)"\times" Quantity of automobiles(2011)

="8\times (200000+20000)\times320"

="1600000+6400000=8000000"


Real GDP 2011=For real GDP we use prices of base year , here 2008 is base year


Real GDP=Price of bread (2008)"\times" quantity of bread (2011)+Price of automobiles (2008)"\times" quantity of automobiles (2011)

"5\times (200000+15000)\times320"


Real GDP in 2011= 1000000+4800000


="5800000"



"Q_b" GDP deflator

="\frac{Nominal GDP}{Real GDP}\times100"


"\frac{8000000}{5800000}\times100=137.93"


This means that Nominal GDP is 37.93% more than the real GDP


"Q_c" inflation rate in 2011


GDP deflator of base year=100

GDP deflator for year 2011=137.93


inflation rate is percentage change in GDP deflator

="137.93-100=37.93"


="\frac{37.93}{100}\times100=37.93 percent"